ra Nat Formula Sheet Previous Quest Free fall means that an object is falling freely with no forces acting upon it except gravity. The distance the object falls, or height, h, is half the product of gravity and the square of the time falling Which of the following represents the equation in terms of t? 3 4 5 6 8 9
Answer #1

h=\frac{1}{2}gt^2

Therefore, the value of t in terms of h is

t^2=\frac{2h}{g} => t=\sqrt{\frac{2h}{g}}

Hence, answer is Option B